### v1

# Understanding Long Videos in One Multimodal Language Model Pass

Abstract
Large Language Models (LLMs), known to contain a strong awareness of world knowledge, have allowed recent approaches to achieve excellent performance on Long-Video Understanding benchmarks, but at high inference costs. In this work, we first propose Likelihood Selection, a simple technique that unlocks faster inference in autoregressive LLMs for multiple-choice tasks common in long-video benchmarks. In addition to faster inference, we discover the resulting models to yield surprisingly good accuracy on long-video tasks, even with no video specific information. Building on this, we inject video-specific object-centric information extracted from off-the-shelf pre-trained models and utilize natural language as a medium for information fusion. Our resulting Multimodal Video Understanding (MVU) framework demonstrates state-of-the-art performance across long-video and fine-grained action recognition benchmarks. Code available at: https://github.com/kahnchana/mvu
Keywords:
Video Understanding Multi-Modal Language Models
1 Introduction
Figure 1: Overview of Framework: We propose three variants of our framework that solves complex long-video question-answering tasks. (left-top) Just LLM utilizes only world knowledge with zero task-specific awareness. (left-right) Single Frame VLM processes an additional center frame to obtain task context but accesses no video specific information. (right) Our complete approach, MVU extracts three additional object-centric modalities followed by fusion in language space. LS refers to likelihood selection.
What can we learn from videos,
beyond scene context understood from a single natural image?
Videos contain unique forms of information beyond what is contained within a single natural image [7]. The recent success of large language models (LLMs) and their visual extensions, Vision-Language Models (VLMs), have resulted in incredible performance on a range of complex language-tied video understanding benchmarks [83, 35, 77, 3], particularly on long-video question answering: a task that requires awareness over longer temporal windows [35] as well as causal and temporal action reasoning [65]. However, the LLMs underlying these approaches contain extensive world knowledge (e.g. understanding of physics, culture, human common-sense [76, 1, 85, 63, 70, 30]), raising the question of whether they excel at video tasks due to actual video modality awareness or simply utilizing world knowledge and contextual information. Such understanding of model reasoning is important for robust deployments avoiding spurious correlation based predictions [80, 49] as well as for model interpretability [51, 25].
We systematically study this question in the context of long-video question-answering benchmarks. First, motivated by the next-token prediction objective in LLM training [45] and the choice selection aspect of long-video benchmarks, we propose an efficient inference strategy utilizing intermediate logits of autoregressive models extracted from a single forward pass, termed likelihood selection. By doing so, we unlock a faster inference ability for autoregressive LLMs/VLMs, that can generate selection results in one forward pass. We then construct two modality-constrained frameworks tagged Just LLM and Single Frame VLM. The first is constrained to access only world knowledge (i.e. no task-specific information beyond the question), whereas the latter has access to task context with an additional single center-frame input. We discover how these models perform significantly better than a random prediction on long-video understanding benchmarks (see Tab. 4). In fact, the latter utilizing purely world knowledge and contextual information outperforms multiple recent state-of-the-art video understanding works (see Tab. 1), challenging the notion of how much video is actually utilized by existing approaches to solve these complex video tasks.
Motivated by these findings, we study how to inject additional video-specific information into our preliminary frameworks using natural language in an interpretable fashion, and whether it can improve their video understanding. We explore three forms of object-centric information modalities, develop pipelines requiring zero video-level training to extract such information, and utilize natural language to fuse this multi-modal information. Our resulting framework, tagged MVU, while achieving state-of-the-art zero-shot performance across long-video understanding benchmarks, also exhibits better interpretability (e.g. exposing video-specific information utilized) through its language-based operation.
In summary, our key contributions are as follows:
  1. 1.
    Likelihood Selection: LLM-based choice selection in a single forward pass
  2. 2.
    Uncovering surprisingly strong performance on complex video-language tasks by resulting modality-constrained baselines
  3. 3.
    MVU framework for language-based fusion of video-specific modalities.
2 Related Work
Video Modality Exploration: Multiple recent works dissect video modality into individual components [80, 7, 50, 48]. Single frame baselines are one interesting sub-class [7, 13, 84, 54, 5]. Extracting object-centric video modalities is another idea, spanning back to [13] which extracts multiple small objects from frames followed by modeling relations across frames and objects. Similarly, [54, 84] combines spatial information with single images to perform video tasks. However, these prior approaches focus on simple video tasks (i.e. action recognition) limited to visual modality. In contrast, our approach tackles the more complex language-tied task of long-video question answering. This task is also explored in [7], but we differ with likelihood selection and multi-modal information fusion.
Long Video Question Answering: Long-video question-answering benchmarks are constructed to specifically test strong causal and temporal reasoning [65] over long temporal windows [35]. Early works explore querying objects or events based on referential and spatial relations [69, 82, 79] followed by focus on temporal modeling of sequential events [27, 28, 20, 66, 67]. More recent works leverage large language models (LLMs) [77, 43, 62, 3, 73, 75] to directly perform these tasks but require video-caption training. In contrast, our MVU operates zero-shot on these tasks requiring no video-level training. Limited zero-shot operation is explored in [61, 83], but we differ in using object-centric modality information and efficient LLM sampling.
Large Language Model Reasoning: Recent LLMs [41, 10, 9] demonstrate multiple forms of strong reasoning abilities [26, 11, 33] including combining different information [64]. Their recent open-source variants [59, 58, 21] achieve equally promising skills using scaled-down models [21] while also demonstrating strong world knowledge [76, 1, 85, 63, 70, 30]. In our work, we leverage these strengths of LLMs for complex video-language tasks, focused on disentangling the effect of their world knowledge from task context and video-specific information.
Language based Fusion: The idea of fusing different modality information using natural language as a medium has been explored in multiple recent works [51, 31, 17, 60, 18, 81]. In [51, 31] language is utilized as an implicit medium for self-supervising video action recognition. Multimodal information represented as language is fused with visual information for action recognition tasks in [17, 60, 18]. We utilize a similar language-as-a-medium fusion of multimodal information but explore this in the context of complex video-language tasks. Work in [81] exploring image and video question-answering tasks is most similar to our work. Our proposed MVU, however, differs with a focus on long-video understanding, object-centric information modalities, and our LLM-specific proposals for efficient selection.
3 Solve in One LLM Pass
In this section, we first present Likelihood Selection, an efficient technique that unlocks autoregressive LLMs to solve multiple selection problems with only one independent forward pass. This is in contrast to next token sampling requiring iterative generations dependent on each previous output. Based on that, in Sec. 4, we introduce Multimodal Video Understanding (MVU), a simple framework that aggregates multimodal video-relevant information using natural language, which achieves significant improvements over baselines across multiple datasets.
3.1 Problem Formulation
We focus on two types of video understanding tasks:
  1. 1.
    Long Video Question Answering (Multiple-Choice-based Selection)
  2. 2.
    Fine-Grained Video Action Recognition (N-Way Classification)
and construct a unified problem formulation for both tasks aforementioned.
Consider a video $x_{v}\in\mathbb{R}^{L\times H\times W\times C}$, a textual question $x_{t}$, a set of textual candidate answers $Y=\left\{y_{i},i=1,...,M\right\}$, and a model $V(\cdot)$ selecting one answer from the given set of answers (noted as $\hat{y}:=V(x_{v},x_{t},Y)$). Selected $\hat{y}$ should ideally be identical to groundtruth $y_{g}$. Here $L,H,W,C$ are the number of frames of the video, frame height, width, and number of channels respectively. $M$ is the number of candidate answers. For multiple-choice-based selection tasks, $x_{v}$, $x_{t}$, and $Y$ are directly present in dataset. For N-Way Classification tasks, we set $x_{t}$ as a generic question (details in Appendix 0.A) and formulate $Y$ by applying a fixed template to the labels of all N classes of the dataset. This formulation is used for the remainder of the paper unless a specific exception is noted.
3.2 Likelihood Selection
Recent autoregressive LLMs and their visual extensions (V-LLMs or VLMs) possess extensive world knowledge that is highly generic and transferable. They are commonly used through iterative prediction of the next tokens conditioned on prior outputs to generate complete natural language outputs. Such a generation process is usually modeled as sampling from a conditional likelihood shown as  Eq. 1, where $\hat{y}^{j}$ stands for the $j^{\mathrm{th}}$ token in a textual sequence $\hat{y}$ autoregressively generated by the model.
$P(\hat{y}|x_{t})$ $=\prod_{j}P(\hat{y}^{j}|\hat{y}^{1,...,j-1},x_{t})$ (1)
The dependency on prior output $\hat{y}^{1,...,j-1}$ makes this process both computationally costly and redundant in the case of choice-based answer selection tasks. Alternately, given the closed set of $Y$ in choice-based selections tasks, we formulate $P(y_{i}|x_{t})$ for any $y_{i}\in Y$ with no dependency on any model generated output ($\hat{y}$) as,
$P(y_{i}|x_{t})$ $=\prod_{j}P(y^{j}_{i}|y^{1,...,j-1}_{i},x_{t})$ (2)
Assume a perfect LLM, intuitively when $y_{i}$ is a proper answer to the question $x_{t}$ (say $y_{i}=y_{g}$), the conditional likelihood $P(y_{i}|x_{t})$ should be larger than any other $P(y_{w}|x_{t})$ where $y_{w}$ is a wrong answer to question $x_{t}$. In fact, modern LLMs are trained with a similar objective [45]. Motivated by this standard LLM training objective, we estimate the relative numerical scales of conditional likelihood on different answers $P(y_{i}|x_{t})$ using a cross-entropy error $e_{i}$, given their equivalence (negative log-likelihood and multiclass cross-entropy, see Section 4.3.4 in [6]). We calculate $e_{i}$ with a single forward pass of LLM without detokenization and the selection can be made by simply picking up the answer with the lowest error, equivalent to the highest conditional likelihood among all the answers.
Figure 2: Likelihood Selection Workflow: We illustrate how our proposed selection strategy can be efficiently parallelized and calculated with a simple cross-entropy loss in one forward pass, followed by an argmin operation, without generating additional tokens. See Sec. 3 for more details.
Therein, we propose Likelihood Selection, first illustrated with a toy example in Fig. 2, where the task is vanilla question-answering with only textual context and the model takes one question $x_{t}$ as well as $M=5$ candidate answers $y_{1,...,5}$. To find the best answer, we simply concatenate the question with each candidate independently ($s_{i}=\mathrm{concat}\left(x_{t},y_{i}\right)$ ) and pad them into a batch $\{s_{1,...,5}\}$. Then the LLM takes the batch of five samples with casual attention masks and performs one inference forward pass, resulting in five shifted logits $\{p_{1,...,5}\}$. Next, we shift the input sequence $s_{i}$ to align the logits $p_{i}$ and calculate the average cross-entropy error only on tokens of $y_{i}$ Finally, the answer with the smallest $e_{i}$ will be picked up as the best answer. The method can be formulated as in Eq. 4 using equivalence of negative log-likelihood to cross-entropy in Eq. 3. Here $n_{i}$ stands for the token sequence length of $y_{i}$ and $p_{i}^{j}$ stands for logits of the $j^{\mathrm{th}}$ token in $p_{i}=V(\mathrm{concat}\left(x_{t},y_{i}\right))$ with logits limited to only those of $y_{i}$.
$e_{i}(y_{i})=\mathrm{CE}(p_{i},y_{i})$ $=\frac{1}{n_{i}}\sum_{j}^{n_{i}}\left(\mathrm{CE}(p_{i}^{j},y_{i}^{j})\right)\approx\sum_{j}^{n_{i}}-\mathrm{log}\>P(y_{i}|x_{t})$ (3)
$\mathcal{F}_{\text{LS}}(Y,x_{t})$ $=\operatorname*{arg\,max}_{y_{i}\in Y}P(y_{i}|x_{t})=\operatorname*{arg\,min}_{y_{i}\in Y}e_{i}(y_{i})$ (4)
In summary, Likelihood Selection performs one single forward pass to extract the network logit outputs, calculates error ($e_{i}$) on each choice, and selects the choice with the lowest error. Note that our method does not utilize any iterative autoregressive generation using the LLM. This results in considerable speed-ups for inference time. We also obtain the additional advantages of avoiding LLM hallucinations and deviations from expected output formats over iterative generation strategies applied to similar visual tasks [18] leading to better accuracy (see Tab. 6.). In Sec. 5.3, we demonstrate both our speed gains and performance improvements.
Furthermore, Likelihood Selection is a generic method that can be easily extended to autoregressive VLMs, and in principle, there is no reason it could not also be used with extra modalities besides language. In the coming section, we will extend this method to a VLM and unlock its ability to process token sequences from multimodalities in one forward pass.
4 Multimodal Video Understanding Framework
We propose a Multimodal Video Understanding Framework that integrates multimodal information from video using natural language. Our approach adapts a pre-trained VLM to construct a powerful long video understanding agent without performing any additional training. We first introduce a VLM-based Frame Selection module, based on a pre-trained VLM and Likelihood Selection, to extract 8 frames per input video, optionally considering the task questions and candidate answers. We then extract three object-centric modalities from the selected frames and use a fixed template to aggregate them into natural language. Finally, by applying likelihood selection again, we successfully turn a pre-trained generative VLM into a discriminator for video-language-based selection tasks. An overview of our framework is illustrated in Fig. 3.
4.1 Frame Selection
Due to the redundancy between video frames, long videos are generally sub-sampled to lower frame rates before processing to solve video tasks. Selecting frames (from a single video) most suitable for each task has been explored under numerous settings [7], with recent works also utilizing LLMs [78]. Motivated by [78], we extend our likelihood selection strategy to this frame selection task, demonstrating the adaptability of our selection strategy for alternate use cases. Specifically, we start from the conditional likelihood of a VLM, noted as $P(y|x_{t},x_{v}^{l})$ where $x_{v}^{l}$ stands for the $l^{\mathrm{th}}$ frame sampled from a video $x_{v}$. We first uniformly downsample the raw video in temporal to $L=16$ frames. For a fixed question $x_{t}$ and target $y$ (details in Appendix 0.B), we iterate through all 16 frames $x_{v}^{l}$ extracted from $x_{v}$ to compare the conditional likelihoods. Finally, we sample the first $8$ images per video in descending order by the conditional likelihood.
Figure 3: Overview of proposed framework for Multimodal Video Understanding, MVU.
4.2 Object-Centric Modalities
Having established a preliminary framework for frame selection and LLM-based answer selection, we replace the LLM with a VLM to process an additional single most relevant frame, $x_{v}^{\tilde{l}}$ ($\tilde{l}$ is the index of the top match frame from frame selection). We then explore how actual video modality information can be leveraged for further improvement. To this end, we focus on object-centric modalities, considering all object categories across the video, spatial locations of individual object instances, and their movement across time. These three forms of information can be extracted from a given video using off-the-shelf pre-trained models.
  1. 1.
    Global Object Information In this stage, we introduce global information that spans beyond a single video frame. For each of the 8 selected frames, we utilize a generative VLM to create object lists and obtain a set of distinct object categories contained across the video.
  2. 2.
    Object Spatial Location Given objects present per frame, we utilize an open-vocabulary object detector to localize each object onto frame coordinates. Additionally, we utilize the similarity of feature vectors (obtained from the object detector) for the same class objects to track object instances across frames. Following prior work [52], we calculate the average center coordinates and scale value for each object instance across all frames. This results in a set of distinct objects $O$ across the video, $O=\{(o_{1},q_{1}),(o_{2},q_{2}),...\}$.  Here, $o_{k}$ describes the object category in natural language while $q_{k}$ contains the x, y coordinates of the object centre and the scale term (area of minimal object bounding box as a ratio to image size, i.e. box area $\div$ image size).
  3. 3.
    Object Motion Trajectory: Next, we leverage the calculated cross-frame object tracks and compute motion trajectories for each object. This modifies our set of distinct objects, pairing each object $o_{k}$ with its trajectory ($o_{k}^{1}\rightarrow o_{k}^{2}\rightarrow...$) across the video frames. We construct an updated set $Z=\{(o_{1},q_{1}^{1}\rightarrow q_{1}^{2}\rightarrow...),(o_{2},q_{2}^{1}\rightarrow q_{2}^{2}\rightarrow...),...\}$. Intuitively, this information should explicitly capture object motion information.
4.3 Language based Fusion
Inspired by [81], we construct our overall framework by injecting these three forms of object-centric information into our setup using natural language. We represent each modality in a fixed template-based fusion. Global object information is represented as a list of category labels, e.g., $x_{\mathrm{GOI}}=$person, oven, dishwasher, …, sink}. Object spatial location modifies this list to include center coordinates ($x,y$) and scale ($s$) where scale is the area percentage occupied by the best-fitting object bounding box. For e.g., $x_{\text{OSL}}=$person located at (0.2, 0.3, 0.07), … , oven located at (0.8, 0.6, 0.04)}. Finally, object motion trajectories update the list to contain frame-level trajectories, e.g., $x_{\text{OMT}}=$person moving as [0.2, 0.3, 0.07] $\rightarrow$ [0.2, 0.4, 0.06] $\rightarrow$ [0.2, 0.6, 0.08], oven moving as … }. Similar to the examples, information from each object-centric modality is represented in textual form to allow their direct fusion and integration into our framework (as additional language inputs). Therein, we describe the resulting setup, our overall framework MVU as follows,
$\hat{y}=\mathcal{F}_{\text{MVU}}(x_{t},x_{v}^{\tilde{l}},x_{\text{GOI}},x_{\text{OSL}},x_{\text{OMT}})$ (5)
We refer the reader to Appendix 0.A for further details on prompting and template operations used as well as examples for our language-based fusion.
4.4 Modality Constrained Variants
We next introduce two modality-constrained variants of our framework tagged Just LLM and Single Frame VLM (see Fig. 1). The former utilizes only the task question injected as a language ($x_{t}$) with no other task-specific information. Note how this variant does not access any information extracted from video. The latter utilizes an additional center visual frame ($x_{v}^{c}$), extracted from the center of the video ($x_{v}$) timeline. This variant accesses no video-specific data (e.g. temporal or motion information). The center frame usage ensures no temporal information leakage in frame selection.
We hypothesize that Just LLM with no access to task-specific knowledge is constrained to generate predictions utilizing its internal world knowledge (e.g. physics, culture, human common sense). We refer to this as world modality. In the case of Single Frame VLM, it has task information but is limited to a single frame, which could possibly provide important scene context. Therein, we refer to this variant as operating in the world and contextual modality information.
Further exploration of these framework variants is presented in Sec. 5.3.
5 Experiments
In this section, we first discuss our experimental setup and datasets. Next, we present our findings when operating in modality-constrained settings, followed by evaluations of our approach on multiple video question-answering benchmarks to demonstrate its performance. Finally, we perform ablations of our framework to verify the usefulness of all proposed components.
Models: Our framework utilizes three distinct off-the-shelf models for its various operations, namely a) an LLM / VLM for likelihood selection (including visual frame selection and textual answer selection), b) a generative VLM for extracting object list from a frame, and c) an open-vocabulary detector for object localization. We use LLaVA-v1.5-13B [32] for likelihood selection, frame selection, and frame object list generation. For object localization, we use OWL-ViT-B/32 [38]. Unless explicitly specified, we use the above setup in all our experiments. Variants of our framework uses LLMs Llama-2-7b-Chat, Gemma-7b-IT, and Mistral-7B-Instruct (default) for likelihood selection. Apart from these off-the-shelf models, our framework involves zero additional training. Also, note that no components of our framework undergo any form of video-level training.
Datasets: We use multiple video datasets for evaluation under question-answering or n-way classification settings. For video question answering, we select two datasets focused on long-form videos: EgoSchema [35], NExT-QA [65]. EgoSchema is a long-form ego-centric video question-answering benchmark, consisting of a 500-video public subset (EgoSchema-S) and a full 5000+ video evaluation set (EgoSchema-F) accessed only through evaluation servers. This dataset spans over 250 hours and is specially constructed to ensure that questions require awareness of a longer temporal window for correctly answering [35]. Example images of EgoSchema are shown in Fig. 4. NExT-QA similarly contains long-form videos with a focus on requiring causal & temporal action reasoning as well as common scene comprehension for correctly answering. It contains a validation set (NextQA-V) of 4996 video-questions pairs and a test set (NextQA-T) of 8564 video-question pairs. We also use a series of robotics datasets from the Open X-Embodiment robotics dataset [40] for video question answering in a different domain (more detail in Sec. 5.2). In our ablations, we utilize a fine-grained action recognition dataset, Something-Something-v2 [15], that contains 174 action categories focusing on object motions by replacing object category nouns with ‘something’ in textual descriptions of each action category.
Compute: Our framework requires no re-training of off-the-shelf models. Our evaluations utilize two NVIDIA RTX A5000 24GB GPUs for inference.
5.1 Long Video Question Answering
Table 1: Ego-Schema Dataset Evaluation: We report top-1 accuracy (%) for video question answering on the open-subset (500 videos) and closed full-set (8564 videos) of Ego-Schema [35]. Our proposed MVU achieves state-of-the-art performance on this benchmark under zero-shot operation with no video level training. We also draw attention to our modality-constrained baselines, Vid-LLM and Vid-VLM, that achieve surprisingly competitive performance on this benchmark. Approaches utilizing proprietary LLMs with significantly higher parameters (e.g. GPT-4) are de-emphasized.
Method
Zero
Shot
Video
Training
Params Frames Subset Full
Random Selection - - - 0 20.0 20.0
ImageViT[43] 1B 16 40.8 30.9
ShortViViT[43] 1B 16 47.9
SeViLA+ShortViViT[43] 5B 32 49.6 31.3
LongViViT[43] 1B 256 56.8 33.3
MC-ViT-L [3] 424M 128 62.6 44.4
FrozenBiLM[73] 1.2B 90 - 26.9
VIOLET[14] 198M 5 - 19.9
mPLUG-Owl[75] 7.2B 5 - 31.1
InternVideo[62] 478M 90 32.1
SeViLA[78] 4B 32 25.7 22.7
LLoVi[83] 13B 180 50.8 33.5
Vamos [61] 13B - - 36.7
LLoVi[83] 1.8T 180 58.3 50.3
Vamos [61] 1.8T - - 48.3
Just LLM (ours) 7B 0 45.8 29.1
Single Frame VLM (ours) 13B 1 55.8 36.4
MVU (ours) 13B 16 60.3 37.6
Long video question answering benchmarks aim to measure the causal and temporal reasoning [65] abilities of models over long temporal windows [35]. In this section, we evaluate our framework on two benchmark datasets and present our results in Table 1 and Table 2.
On EgoSchema dataset, results reported in Tab. 1 demonstrate the state-of-the-art performance of our framework in comparison to work operating under similar settings. We reiterate how our approach is both zero-shot and requires no video-level training. In contrast, multiple prior zero-shot works utilize pseudo-labels on the downstream task videos for transductive training [14, 75, 78]. We also highlight how our MVU even outperforms multiple recent supervised (video-caption) approaches [43]. Among zero-shot methods, while our results are subpar to the trillion parameter variants of [83, 61] that utilize proprietary closed language models, we outperform with a clear margin against the fairer comparison to their similar parameter variants (more in Appendix 0.D). We also draw attention to our two modality-constrained model variants, Just LLM and Single Frame VLM that achieve surprisingly competitive results on this task, even outperforming multiple recent works. Motivated by these results, we conduct further experimentation under modality constrained setting in Sec. 5.3.
Table 2: Next-QA Dataset Evaluation: We report top-1 accuracy (%) for the test-split of Next-QA dataset [65]. Our proposed MVU achieves state-of-the-art results under zero-shot settings with no video-level training. We also highlight our competitive performance across the dataset’s causal, why, and how sub-categories. In table header, ZS stands for zero-shot and VT stands for video level training.
Method ZS VT Params Causal Why How All
Random Selection - - 20.0 20.0 20.0 20.0
CoVGT [68] 149M 58.8 57.4 69.3 60.0
SeViT [23] 215M - - - 60.6
HiTeA [74] 297M 62.4 58.3 75.6 63.1
InternVideo [62] 478M 62.5 58.5 75.8 63.2
MC-ViT-L [3] 424M - - - 65.0
BLIP-2 [29] 4B 70.1 65.2 80.1 70.1
SeViLA [78] 4B 74.2 69.4 81.3 73.8
LLama-VQA-7B [24] 7B 72.7 69.2 75.8 72.0
Vamos [61] 7B 72.6 69.6 78.0 72.5
Just-Ask [72] 66M 31.8 30.4 36.0 38.4
VFC [39] 164M 45.4 51.6 64.1 51.5
InternVideo [62] 478M 43.4 48.0 65.1 49.1
SeViLA[78] 4B 61.3 61.5 75.6 63.6
CaKE-LM w HGA [57] 2.7B 35.3 34.7 36.9 34.8
CaKE-LM w CoMem [57] 2.7B 35.7 35.3 36.8 34.9
LLoVi [83] 13B 55.6 47.9 63.2 54.3
MVU (ours) 13B 55.4 48.1 64.1 55.2
We next evaluate our framework on the NextQA benchmark and report these results in Tab. 2. Our MVU achieves competitive results under zero-shot settings. While [78] outperforms our approach, we note how they require video-caption localization pretraining and appears to overfit to this dataset considering their relatively lower performance on other datasets (see Tab. 1). Considering the zero-shot, video training-free aspect of our proposed MVU, we achieve state-of-the-art results in comparison to similar prior work.
Table 3: Robotics Domain Action Recognition Evaluation: We report top-1 accuracy (%) for our VideoQA style action recognition formulation (details in Sec. 5.2) on the Open X-Embodiment collection of datasets [40]. Our proposed MVU achieves clear improvements over random selection and a single-frame baseline. In the table header, Obs. (observation), size, CC (class count) stand for the camera used, number of videos, and number of unique language instructions per dataset, respectively. In the observation column, T stands for third-person view, a stationary camera that does not move with the robot. While F means first-person view where the camera is mounted on the moving robot. SF-VLM stands for Single Frame VLM.
Dataset Obs. Size CC Random SF-VLM MVU
ASU TableTop Manipulation [87, 86] T 110 83 13.6 20.9 20.9
Berkeley MVP Data [47] F 480 6 20.0 22.9 33.1
Berkeley RPT Data [46] F 908 4 24.6 20.6 20.6
CMU Play Fusion [8] T 576 44 20.3 34.2 35.6
CMU Stretch [2, 37] T 135 5 23.0 18.5 24.4
Furniture Bench [19] T 5100 9 20.2 12.1 16.1
Furniture Bench [19] F 5100 9 20.2 13.1 12.7
CMU Franka Pick-Insert Data [55] T 631 7 18.7 18.9 21.2
CMU Franka Pick-Insert Data [55] F 631 7 23.1 59.6 49.3
Imperial F Cam T 170 17 20.0 24.7 24.1
Imperial F Cam F 170 17 23.5 23.5 20.6
USC Jaco Play [12] T 1085 89 21.8 16.2 18.0
USC Jaco Play [12] F 1085 89 19.4 28.0 32.4
NYU ROT [16] T 14 12 21.4 50.0 57.1
Roboturk [34] T 1959 3 34.7 43.1 37.9
Stanford HYDRA [4] T 570 3 35.1 44.6 68.2
Stanford HYDRA [4] F 570 3 31.2 44.0 48.9
Freiburg Franka Play [53, 36] F 3603 406 20.4 32.2 31.6
Freiburg Franka Play [53, 36] T 3603 406 19.7 21.8 24.0
LSMO Dataset [42] T 50 2 34.0 50.0 72.0
UCSD Kitchen [71] T 150 8 19.3 32.0 32.7
Austin VIOLA [88] T 150 3 26.7 33.3 33.3
Austin VIOLA [88] F 150 3 30.0 33.3 33.3
Total - 27000 - 22.1 23.5 25.0
5.2 Robotics Domain Action Recognition
In this section, we evaluate the generalization capabilities of the proposed methodologies across various subsets of the Open X-Embodiment robotics dataset [40]. The Open X-Embodiment dataset is an extensive collection of visuomotor robotics datasets, encompassing a wide range of tasks and environments. It is designed to facilitate research in visuomotor control, providing rich sensory inputs and motor outputs for training and testing robotic systems. However, the videos are usually taken in a controlled environment and they do not always contain meaningful objects, which makes the samples in the dataset usually out of general video distribution (See Fig. 4).
For our analysis, we specifically select subsets of the dataset that contain expert episodes accompanied by corresponding language instructions and adapt them into video classification datasets. We treat each trajectory within the dataset as a video clip, with its associated language instruction serving as the video caption (classification label). For each episode, the model is tasked with identifying the correct language instruction from a set of five options, considering a video clip uniformly downsampled to 8 frames. The incorrect options are randomly chosen from the dataset to ensure a diverse and challenging selection. In instances where the datasets have multiple cameras for each episode, we treat the videos captured by each camera as distinct datasets.
Figure 4: Examples frames of different videos from EgoSchema dataset (top) and examples from Open X-Embodiment dataset (bottom). Compared to top row, images below appear out of distribution given the controlled environment and robot movement.
Our results presented in Tab. 3 indicate clear performance improvements for our proposed MVU over both random selection and our competitive single frame baseline. The purpose of this experiment is to evaluate the generality of our approach to video domains different from everyday natural videos. We take these promising results in Tab. 3 to indicate the strong generality of our framework.
5.3 Modality Constrained Evaluation
We evaluate the two modality-constrained variants of our approach, Just LLM and Single Frame VLM (details in Sec. 4.4) and summarize these findings in Tab. 4. We uncover surprisingly strong performance of both variants on two long-video understanding benchmarks. Note how these approaches use no video-specific information to generate predictions.
We highlight how our best Just LLM variant achieves performance significantly higher than random selection (+25.8% on EgoSchema-S / +20.1% on NextQA-T) using zero visual information. This indicates the large portion of questions in existing video-QA benchmarks that can be answered correctly purely using world knowledge. We also highlight our single frame variant performing on par with some existing state-of-the-art (gray). In particular, for EgoSchema-S we outperform [83] which uses information extracted from 180 frames per video incurring an inference cost over 100 times higher than ours. In light of these findings, we argue that long video understanding approaches in particular must focus on learning information beyond what a single frame baseline can achieve.
We also take these results as an indication of the clear inference speed-up from Likelihood Selection (more specific ablations in Sec. 5.4).
Table 4: Modality Constrained Long Video Evaluation: We evaluate our proposed Vid-LLM and Vid-VLM setup, on two video question-answering datasets that specifically focus on longer videos. Number of samples in each dataset inside brackets. EgoSchema-S is public subset of EgoSchema [35] and NextQA-T is the test set of NextQA dataset [65]. LS backbone is the pre-trained model used for likelihood selection. We report Top-1 Accuracy (standard evaluation protocol) and inference time per sample (single video-question pair) when running on two NVIDIA RTX A5000 24GB GPUs. Recent state-of-the-art work [83] using multi-frame inputs (gray) is also reported for comparison. indicates the use of additional LLMs/VLMs for intermediate processing.
Method LS Backbone EgoSchema-S (500) NextQA-T (8564)
Acc (%) Time (s) Acc (%) Time (s)
Random - 20.0 - 20.0 -
Just LLM (ours) Llama-2-7b-Chat 17.4 0.89 28.8 1.17
Just LLM (ours) Gemma-7b-IT 45.8 2.38 39.3 2.60
Just LLM (ours) Mistral-7B-Instruct 45.8 0.41 40.1 0.55
SF-VLM (our) LLaVA-v1.5-13B 55.8 1.89 51.2 2.03
LLoVi [83] Mistral-7B-Instruct 50.8 381 54.3 207
5.4 Ablations
We systematically dissect our overall framework to establish the usefulness of each individual component of our method. We first ablate our overall MVU framework and report these results in Tab. 5. Our evaluations indicate clear performance boosts from both our proposed frame selection strategy and each of our object-centric modality information.
We next perform ablations on our likelihood selection strategy and report these results in Tab. 6. Our results indicate clear performance boosts due to likelihood selection (LS) in terms of both accuracy and inference speed. In the alternate case (no LS), a standard post-generation selection [83] (i.e. generate an answer and match against ground-truth selection) is utilized. While our inference speed boosts against this setup can be directly attributed to our single pass aspect, we hypothesize that accuracy gains arise as a result of reduced LLM output hallucination and deviations from expected output formats, that are commonly observed in other approaches [18].
We also conduct an ablation on our object motion trajectory modality inputs focussed on the fine-grained action recognition dataset, Something-Something-v2 [15] (SSv2). As elaborated in Sec. 5, SSv2 is a dataset focused on motion-based category discrimination, providing an ideal evaluation to measure the usefulness of our object motion trajectory modality. We benchmark on a subset of this dataset following [31] and report these results in Tab. 7. Our results while exceeding their performance also indicate the clear performance gains obtained when injecting the object motion trajectory modality into our MVU framework.
Table 5: Ablation on object-centric modality information. We report accuracy (top-1 %) on the 500 sample subset of EgoSchema. In the table header, VI and FS stand for visual inputs and frame selection respectively. GOI, OSL, OMT refer to our object-centric modalities (see Sec. 4.2). Each component of our approach, FS and the integration of separate modality information, results in clear performance improvements. Our overall framework, MVU (our-full), achieves the best performance.
Method VI FS GOI OSL OMT EgoSchema-S
Just LLM (ours) 45.8
Single Frame VLM (ours) 55.8
MVU (ours) 56.4
MVU (ours) 58.6
MVU (ours) 58.8
MVU (ours-full) 60.3
Table 6: Ablation on Likelihood Selection. Results indicate performance improvements in terms of both top-1 accuracy (%) and inference time (s) when utilizing our likelihood selection approach.
Method LS Accuracy Time
MVU (ours) 56.4 8.65
MVU (ours) 60.3 2.42
Table 7: Ablation on OMT modality. Results in the motion-based SSv2 dataset indicate the usefulness of our OMT modality.
Method OMT Accuracy
Random - 0.6
CLIP [44] - 4.0
MAXI [31] - 6.4
MVU (ours) 3.6
MVU (ours) 7.2
6 Conclusion
In this work, we present a multimodal video understanding framework, termed MVU, that achieves state-of-the-art performance on complex video understanding tasks. In particular, the evaluation of long-video question answering and fine-grained action recognition (language-tied zero-shot) demonstrates the strong performance of our MVU framework. We propose an efficient technique for LLM-based answer choice selection and extend this for frame selection highlighting its generality. We separate video-specific information into three object-centric modalities, propose a language-based fusion of this multimodal information, and explore their individual contributions to our end performance.
Revisiting our original motivation on, “what we can learn from videos, beyond scene context understood from a single natural image”, in this work we also presented two modality-constrained variants of our approach that uncover surprising insights relevant to this question. We achieve strong results on long-video understanding benchmarks using no video-specific data, and build over that baseline to showcase the additional performance gains achievable through injecting video-specific information.
Reproducibility Statement
Our method utilizes multiple pretrained models, all of which are open-source with model weights freely available. We use the versions of these models hosted on HuggingFace https://huggingface.co/ for all our experiments. We discuss all steps in our proposed algorithms in detail while also releasing relevant code. All evaluations we perform are on public datasets accessible by all (some behind evaluation servers to prevent test set contamination). Our code is hosted at https://github.com/kahnchana/mvu.
Ethics Statement
We present a framework specific for long-video understanding fusing individual video modalities through language and also explore efficient inference strategies for LLM-based methods. Although we believe our algorithms and framework introduce no additional negative biases (especially given how we do not train any new models), we acknowledge our use of existing models that are known to contain certain biases (from the datasets they are trained on). However, in our work, we consider some of these biases (i.e. world knowledge) as actual strengths that should be given careful consideration when evaluating such LLM-based frameworks even on complex downstream tasks. Additionally, we highlight the interpretability of our language-based information fusion, which allows explicit observation of the types of biases present, allowing easier alterations for cases that require the absence of certain biases.
Acknowledgements
This work was supported by the SUNY Empire Innovation Program (EIP). We thank the Robotics Lab at SBU, especially Jongwoo Park, for helpful discussions.
Contributions
KR led the project coming up with the initial idea, building the preliminary codebase, and performing most experimentation. XL was responsible for extending the framework to the robotics domain, contributed ideas for extracting OSL and OMT modality information, and constructed all figures in the paper. KK ran preliminary experiments on verifying our likelihood selection strategy, performed multiple long-video question-answering experiments, and discussed all aspects of the project. MR organized and supervised the entire project, set the research direction, and discussed all aspects of the project.
Appendix
Appendix 0.A Prompting and Template Operations
In Sec. 4.2 and Sec. 4.3, we utilize 3 distinct prompts and fusion templates for generating joint textual inputs to be processed by the LLM. The 3 distinct prompt categories correspond to our Global Object Information ($x_{\mathrm{GOI}}$), Object Spatial Location ($x_{\mathrm{OSL}}$), and Object Motion Trajectory ($x_{\mathrm{OMT}}$) modalities. We first describe our exact templates as Python pseudo-code in Table A.1.
[…]
Table A.1: Prompt templates for three textual modalities.
The above templates depend on each of the modalities represented in textual form (i.e. GOI_data, OSL_data, OMT_data). We describe their exact textual forms next using examples in Table A.2.
GOI_data = ["person", "oven", "dishwasher", "sink", "countertop", "dish", "box", "scissors", "drain", "hand", "stove"]   OSL_data = ["stove located at (0.52, 0.64, 0.595)",       "sink located at (0.56, 0.64, 0.211)",       "countertop located at (0.63, 0.79, 0.308)",       "box located at (0.46, 0.65, 0.142)",       "dishwasher located at (0.5, 0.5, 0.991)",       "dish located at (0.41, 0.75, 0.077)",       "person located at (0.47, 0.76, 0.282)" ]   OMT_data = ["stove trajectory: (0.5,0.5,0.991)->(0.51,0.69,0.397)       ->(0.54,0.73,0.396)",       dish trajectory: (0.55,0.62,0.096)->(0.11,0.65,0.079)",                   .                   .                   .       "dish trajectory: (0.41, 0.75, 0.077)",       "person trajectory: (0.54,0.81,0.34)->(0.49,0.72,0.339)        ->(0.54,0.84,0.157)->(0.23,0.71,0.176)        ->(0.51,0.79,0.232)->(0.52,0.78,0.266)        ->(0.39,0.64,0.558)->(0.54,0.82,0.184)"]
Table A.2: Prompt examples for three textual modalities.
[…]
Appendix 0.B Frame Selection Queries
As described in Sec. 4.1, we utilize our likelihood selection strategy for selecting frames optimal to solve a selected task. In this setting, we fix the textual input questions $x_{t}$ and possible target $y$ (only one target) and alter the input visual tokens based on different frames $x_{v}^{l}$ ($l$ is the index of video frame), in order to obtain per-frame likelihood measures. The prompt template at presented at Table A.3
User <image to select $x_{v}^{l}$> Consider task: task_question. How useful is this frame to answer the given question?   Assistant This frame is highly useful and contains relevant information
Table A.3: Prompt template for frame selection.
We hypothesize that frames with more relevant information will maximize the likelihood of our target text and utilize our promising results (see performance gains from frame selection in Table 5) as validation.
Appendix 0.C Tokenization in LLMs
Most modern LLMs utilize Byte-Pair Encoding (BPE) tokenization [56] to convert natural language into discrete tokens that are mapped to vocabulary embeddings. This process is learned from language itself and the resulting tokenization may sometimes break complete words into pieces (e.g. example $\rightarrow$ ex-am-ple). Given our utilization of logits extracted from an LLM forward pass, we note that each logit corresponds to a single token, which may at times be the embedding of some meaningless word piece. However, our calculation of a joint likelihood across a sequence of tokens ensures a meaningful process, which is validated by our strong results.
Appendix 0.D Details on Baselines
In Sec. 5.1, we evaluate the performance on long-video understanding tasks using work in [83, 61] are two baselines for comparison. However, both these methods utilize closed-source, proprietary LLMs (i.e. GPT-4) with parameter counts on the order of trillions (over 100X our model size) deeming their unfair comparisons. In the case of [83], we replicate their method (using their open-source repository and pre-trained models following [22]) utilizing an open-source LLM of comparable parameter count as our framework. For [61], the authors directly report results for a variant with a similar parameter count as ours. We utilize these evaluations as our point of comparison.